OpenStudy (anonymous):

When 5 is added to 2 times a number, the result is 45. Find the number. A. 20 B. 45 C. 5 D. 2

OpenStudy (texaschic101):

5 + 2x = 45 2x = 45 - 5 2x = 40 x = 40/2 x = 20

OpenStudy (anonymous):

45-5 = 40 40/2 = 20 20 is the answer